An automolbile buyer is trying to decide whether to buy a new car and finance it for 5 years at 7% interest or a used car costing $11500 less at 11% for 2 years. The new car load will cost $5130 more in interest than the used car loan. Find the cost of each car
Answer by mananth(16949)

An automolbile buyer is trying to decide whether to buy a new car and finance it for 5 years at 7% interest or a used car costing $11500 less at 11% for 2 years. The new car load will cost $5130 more in interest than the used car loan. Find the cost of each car
...
let new car price be $x
interest on this car = 0.07*x*5=0.35x
...
used car
price = x-11500
interest = 0.11*(x-11500)**2
..
0.35x - 0.11*(x-11500)*2=5130
0.35x-0.22x+2530=5130
0.13x=2600
x=2600/0.13
x=$ 20,000 new car
..
old car = $,9500
...
m.ananth@hotmail.ca
.